OpenStudy (anonymous):

What's the laplace transform of 2t? I thought it was 2/s, but now I am not certain.

OpenStudy (abb0t):

Can you show your steps of how you got \(\large \frac{ 2 }{ s }\)?

hartnn (hartnn):

or do you know the general formula for Laplace transform of t^n ?

OpenStudy (anonymous):

The transform table says that L{t^n} = n!/(s^(n+1))

hartnn (hartnn):

yes, so to find L.T of 't' you'll put n=1 right ? (t=t^1) what do you get ?

OpenStudy (anonymous):

I got (1/s^2) * 2

hartnn (hartnn):

thats correct L.T for 2t = 2/s^2

OpenStudy (anonymous):

Thanks